# Adversarially-Robust Gossip Algorithms for Approximate Quantile and Mean Computations

### v2

@@ 2 Approximate Median
where the last inequality holds since $K=\lceil 8\cdot\log_{20\gamma^{\prime}}\left(\frac{\gamma}{4}\right)\rceil$. Notice that $K>0$ because $120\gamma^{\prime}\leq\frac{20}{27}<$ 1. Thus, the expected number of nodes that do not output a correct approximate median is at most $\frac{\gamma^{4}n}{256}$.
@@
3 Shifting Quantiles
In this section, we prove the theoretical runtime guarantees of Theorem 2 for our robust quantile-shifting algorithm. See 2 Note that the assumptions $\beta\leq\frac{\varepsilon^{2.5}}{16}$ and $\varepsilon\leq\frac{1}{6}$ together imply that $\beta\leq\frac{\varepsilon\cdot\varepsilon^{3/2}}{16}\leq\frac{\varepsilon}{16}\cdot\frac{1}{6\sqrt{6}}=\frac{\varepsilon}{96\sqrt{6}}\leq\frac{\varepsilon}{200}$. Without loss of generality, we will also assume that $\phi\leq 1/2$, as the algorithm and analysis is exactly symmetric for the other case.
The pseudocode for our algorithm is described in Algorithm 2. Let $L_{i},M_{i}$ and $H_{i}$ denote the nodes whose quantiles lie in $\left[0,\phi-\varepsilon\right),\left[\phi-\varepsilon,\phi+\varepsilon\right]$ and $\left(\phi+\varepsilon,1\right]$ respectively, at the end of iteration $i$ of the while-loop. We want to show that $|L_{i}|$ and $|H_{i}|$ decrease rapidly as the algorithm progresses.
$h_{0}\coloneqq h_{0}^{\prime},\qquad h_{i+1}\coloneqq(h_{i}+\beta)^{2},$
Proof Sketch. Let $t$ be the total number of iterations of the while-loop executed by Algorithm 2. We first show that after $t$ iterations, $h_{t},h^{\prime}_{t}$ are (still) close to each other. Next, we show that for all $0\leq i\leq t$, $|H_{i}|/n$ is bounded above by $h^{\prime}_{t}$ in expectation and by $h_{t}$ with high probability. Thus, at the end of the algorithm, $|H_{t}|/n$ is close to (both) $h_{t}$ and $h^{\prime}_{t}$, which are both close to $\frac{1}{2}-\varepsilon$. We also show that $|M_{t}|/n\geq 2\varepsilon$ w.h.p. These two statements put together will then imply our theorem.
We begin with a deterministic runtime bound for Algorithm 2.
@@ 3 Shifting Quantiles
Note that lines $9,10$ in Algorithm 2 are executed with probability $1$ during the first $t-1$ iterations. Indeed, $\delta$ is set to $1$ if and only if $h^{\prime}_{i}-h^{\prime}_{i+1}\leq h^{\prime}_{i}-T$ or equivalently $T\leq h^{\prime}_{i+1}$, which is true up to $i=t-1$ by definition of $t$. We now prove that $h_{t},h^{\prime}_{t}$ are close to each other, which we will later use to show that $H_{t}$ is concentrated around its expectation.
@@
For the upper bound on $h_{t-1}$, notice that by Claim 13, $h_{t-1}\geq h^{\prime}_{t-1}\geq 9/32$. Now, we will show that for all $20\leq i\leq t-$ 2, we have $h^{\prime}_{i}>10/19$. Suppose that by contradiction $h^{\prime}_{i}\leq 10/19$ for some such $i$. This must mean that $h^{\prime}_{i+1}\coloneqq\left(h^{\prime}_{i}-\beta\right)^{2}\leq\left(\frac{10}{19}\right)^{2}<\frac{9}{32}$, which is a contradiction. Thus, we have $h_{i}\geq h^{\prime}_{i}\geq 10/19$ for all $20\leq i\leq t-$ 2. Using this and applying the bounds on $h_{t-1}$ and $h^{\prime}_{t-1}$ iteratively, we get
@@
For iteration $11\leq i\leq t-$ 1, $\left(\frac{|H_{i}|}{n}-\beta\right)^{2}\leq\mathbb{E}\left[\frac{|H_{i+1}|}{n}\mid H_{i}\right]\leq\left(\frac{|H_{i}|}{n}+\beta\right)^{2}$
@@
Suppose that $(1-\varepsilon^{\prime\prime})h^{\prime}_{t-1}\leq\frac{|H_{t-1}|}{n}\leq(1+\varepsilon^{\prime\prime})h_{t-1}$ for some $916\beta\leq\varepsilon^{\prime\prime}\leq 1/$ 9. Then we have $T-\frac{3}{2}\cdot\varepsilon^{\prime\prime}\leq\mathbb{E}\left[\frac{|H_{t}|}{n}\mid H_{t-1}\right]\leq T+\frac{3}{2}\cdot\varepsilon^{\prime\prime}+\frac{1}{2}\cdot\varepsilon$.
@@
In particular $\mathbb{E}\left[\frac{|H_{t}|}{n}\mid H_{t-1}\right]=\Theta(1)$ (i.e. $\mathbb{E}\left[\frac{|H_{t}|}{n}\mid H_{t-1}\right]$ is bounded away from 00). Thus, by using Chernoff bounds with $\varepsilon^{\prime}$, we get w.h.p.
@@
In the rest of this section, we aim to show that $M_{i}$ does not change much over the course of the algorithm. First, we investigate the change in the value of $|M_{i}|$ for the first $t-1$ iterations of the algorithm.
@@
For $10\leq i<t-$ 1, suppose that $\frac{|H_{i}|}{n}\geq\frac{1}{2}$ and $\frac{|M_{i}|}{n}\geq\frac{|M_{0}|}{n}$, then $\mathbb{E}\left[\frac{|M_{i+1}|}{n}\mid M_{i},H_{i}\right]\geq\frac{|M_{i}|}{n}\cdot\left(1+\frac{7\varepsilon}{4}\right)$.
@@
For $10\leq i\leq t-$ 1, suppose that $\frac{|M_{i}|}{n}\geq\frac{|M_{0}|}{n}$, then
@@
First, note that for $20\leq i\leq t-$ 2, w.h.p., we must have $\frac{|H_{i}|}{n}>1/2$. Otherwise, by Lemma 15 and using Chernoff bounds with error $\varepsilon^{\prime}=1/100$, and the inequalities $\varepsilon\leq\frac{1}{6},\beta\leq\frac{\varepsilon}{200}\leq\frac{1}{100}$, we have w.h.p.
@@
Next, we show by induction that w.h.p. $\frac{|M_{i}|}{n}\geq\frac{|M_{0}|}{n}$ for $10\leq i\leq t-$ 1. The claim is trivially true for $i=0$. Otherwise, suppose that $\frac{|M_{i}|}{n}\geq\frac{|M_{0}|}{n}$ is true for some $i$. Using Lemma 19 and 21 with $\varepsilon^{\prime}\coloneqq\varepsilon/4$, we have
@@
By Lemmas 18 and 22, we have $\frac{1}{2}-\frac{15\varepsilon}{8}\leq\frac{|H_{t}|}{n}\leq\frac{1}{2}-\frac{\varepsilon}{8}$ and $\frac{|M_{t}|}{n}\geq 2\varepsilon$ w.h.p. Consequently, we must have $\frac{|H_{t}|}{n}+\frac{|M_{t}|}{n}\geq\frac{1}{2}+\frac{\varepsilon}{8}$. Thus, we conclude that any $\phi^{\prime}$-quantile where $\phi^{\prime}\in\left[\frac{1}{2}-\frac{\varepsilon}{8},\frac{1}{2}+\frac{\varepsilon}{8}\right]$ must be in $M_{t}$. ∎
4 Approximate Mean
The objective of this section is to prove the runtime and correctness guarantees for our adversarially robust approximate mean algorithm. We recall Theorem 3 here for the convenience of the reader.
@@ 4 Approximate Mean
The pseudocode is described in Algorithm 3. In the first phase of the algorithm, in each round, every node $u$ samples two nodes and updates its value to their arithmetic mean. Sampled node values that exceed the interval of acceptable values $[0,M]$ are clipped to the interval boundaries. In the second phase, the algorithm globally samples a set of nodes uniformly at random and outputs the median value of these nodes. Let $x_{v}(t)$ denote the value $x_{v}$ stored in the node $v$ at the end of iteration $t$. Additionally, let us denote by $V$ the set of all nodes.
To help analyze the progress of the algorithm, we define the potential function
$\Phi(t)\coloneqq\sum_{u\neq v\in V}(x_{u}(t)-x_{v}(t))^{2},$
where the sum goes over all subsets of $V$ of size $2$ (i.e. we do not double-count the pairs). Notice that this potential captures the variation in the values stored in the nodes. As the algorithm progresses, we expect the potential to decrease to a small value, as the values stored in the nodes converge to the average.
We remark that by assumption $\beta\leq\left(\frac{\varepsilon}{100}\right)^{2.5}$ and $\delta<<\left(\frac{\varepsilon}{100}\right)^{2.5}$, which implies that $\eta\leq\left(\frac{\varepsilon}{100}\right)^{2.5}$. Furthermore, we can rewrite $\eta$ as
@@
We also define the quantity $\psi(t)\coloneqq\sum_{u\in V}x_{u}(t)$ to be the sum of the node values. Notice that $\frac{\psi(0)}{n}$ is the true average that we want to approximate.
Proof Sketch. We start by showing that $\Phi(t)$ is small at the end of the first phase of Algorithm 3, and hence the vast majority of the node values are close to each other at that point. Additionally, we show that $\psi(t)$ does not change much throughout the algorithm, and hence the average of all the nodes values stays close to the true average. Hence, the vast majority of the nodes have a correct approximate mean at this point. Finally, after the second phase, we show that the number of nodes with an incorrect value is less than $\gamma n$.
We first give two alternate expressions for $\Phi(t)$ in terms of $\psi(t)$.
@@
Now, notice that $\Phi(t+1)$ is a function of $2n$ independent random variables that correspond to the $2$ nodes chosen by each of the $n$ nodes in this round. Each of these variables appear in exactly $n-1$ of the summation terms and can affect this term by at most $M^{2}$. Thus, the effect of each of these variables on $\Phi(t+1)$ is bounded by $nM^{2}$, so the sum of squares of the effects of all these random variables is bounded by $2n^{3}M^{4}$.
@@
5 Lower Bounds
In this section, we prove Theorem 4, which yields a general tool to derive lower bounds for the round complexity of gossip algorithms in the presence of adversaries. In essence, the theorem quantifies a lower bound on the number of rounds needed for all but a $\gamma$ fraction of the nodes to output a correct answer, in the presence of $\beta n$ adversaries. Note that a similar lower bound was proven for the non-adversarial setting in [HMS18], but its proof does not extend readily to our case because our parametrized notion of correctness allows for a $\gamma$ fraction of nodes to be incorrect in addition to the $\beta n$ adversarial nodes. What remains true however is that even with the slack provided by $\gamma$, a correct algorithm cannot terminate in fewer than $\Omega\big{(}\frac{\log(1/\gamma)}{\log(1/\beta)}\big{)}$ rounds, since up to this point, more than a $\gamma$ fraction of nodes will have communicated only with bad nodes. Therefore they will have received no accurate information whatsoever and could not possibly output the correct result. This is formalized in Theorem 4.
See 4
Proof.
First, suppose that $2\gamma>\beta$. Then, notice that $\frac{\log(1/2\gamma)}{\log(1/\beta)}<1$ and hence the claim holds trivially. Thus, we can henceforth assume that $2\gamma\leq\beta$.
Consider a single good node $v$. In one round, the probability that it communicates with an adversarially affected node is $\beta$. Thus, after $t<\frac{\log(1/2\gamma)}{\log(1/\beta)}$ rounds, the probability that it has only communicated with adversarial nodes is at least $\left(\beta\right)^{t}>2\gamma$. Thus, in expectation after $t$ rounds there is at least a $2\gamma$ fraction of such nodes.
Now, notice that the communication partner chosen by an arbitrary node in an arbitrary round is completely independent of what is chosen by any other node in any round. Thus, by Chernoff bounds, the probability that there are at least a $\gamma$ fraction of the nodes that have only communicated with bad nodes during the first $t<\frac{\log(1/2\gamma)}{\log(1/\beta)}$ rounds is at least
Similarly, one can adapt the proof of [HMS18, Theorem 1.3] to show a $\Omega\left(\log\log\frac{1}{\gamma}+\log\frac{1}{\varepsilon}\right)$ lower bound for these problems. We present the full proof for completeness.
Proposition 31.
@@ 5 Lower Bounds
Together, Theorem 4 and Proposition 31 show that our algorithm for the approximate median is optimal. Moreover, our algorithm for the approximate mean is also almost optimal with only an extra $O(\log\frac{1}{\gamma+\beta})$ factor, and is optimal whenever $\varepsilon$ is polynomially smaller than either $\gamma$ or $\beta$.
Appendix A Appendix
In this appendix, we collect useful tools and bounds that we use for our analysis throughout the paper. We begin with a statement of the standard Chernoff bounds.
Lemma 32.
The following simpler statements hold for $10<\delta<$ 1
Lemma 33.
Suppose that $X_{1},X_{2},\ldots,X_{n}$ are independent indicator random variables, and let $X=\sum_{i=1}^{n}X_{i}$ be their sum. Then, for some $M\geq\mathbb{E}\left[X\right]$ and $10<\delta<$ 1,
Proof.
Lemma 32 and [HMS18, Lemma A.2] immediately imply the desired claim. ∎
Next, we state Azuma’s inequality.
Lemma 34.
$b_{n,k}(p)=\binom{n}{k}p^{k}(1-p)^{n-k},$
$B_{n,k}(p)=\sum_{0\leq l\leq k}b_{n,l}(p).$
Lemma 35.
Proof.
We have
Thus,
Lemma 36.
Proof.
The claim is clearly true for $\alpha=0$. For $\alpha>0$, note that
$(1-2\alpha)^{n}=1-f(n,\alpha)\cdot\alpha$
where
Next, we prove another upper bound for exponentials.
Lemma 37.
Suppose that $0\leq x\leq M$. Then we have
$e^{\alpha x}\leq 1+\frac{e^{M\alpha}-1}{M}x.$
Proof.
The claim is clearly true for $x=0$. For $x>0$, note that
$e^{\alpha x}=1+f(x)x$
where
$f(x)\coloneqq\frac{e^{\alpha x}-1}{x}$
is monotonically increasing in $x$ in its domain. Thus, we have
$e^{\alpha x}\leq 1+f(M)x\leq 1+\frac{e^{M\alpha}-1}{M}x.$
Lemma 38.
Consider the functions $f(x)\coloneqq\left(\frac{x}{1+\beta}\right)^{2}$ and $g(x)\coloneqq\left(\frac{x+\beta}{1+\beta}\right)^{2}$ for some $10<\beta<$ 1. Then, for all $\beta^{2}\leq x\leq 1$, we have $f(x)\leq x$ and $g(x)\leq x$.
Proof.
First, consider the function $F(x)\coloneqq f(x)-x$. We have $F(0)=f(0)-0=0$, and $F((1+\beta)^{2})=f((1+\beta)^{2})-(1+\beta)^{2}=0$. Thus, since $F$ is a quadratic in $x$ with a positive leading coefficient, we must have $F(x)\leq 0$ for all $0\leq x\leq(1+\beta)^{2}$. Consequently, $f(x)\leq x$ for all $\beta^{2}\leq x\leq 1$.
Next, consider the function $G(x)\coloneqq g(x)-x$. We have $G(1)=g(1)-1=0$, and $G(\beta^{2})=g(\beta^{2})-\beta^{2}=0$. Thus, since $G$ is a quadratic in $x$ with a positive leading coefficient, we must have $G(x)\leq 0$ for all $\beta^{2}\leq x\leq 1$. Consequently, $g(x)\leq x$ for all $\beta^{2}\leq x\leq 1$. ∎
Theorem 39.
$\overline{B}_{(2k+1),k}(p)\leq\frac{1}{2}-\frac{13}{12}\cdot\alpha.$
Proof.
For $l>k+1$, notice that $12l-2k-1>$ 1, and hence we can use Lemma 36 to further bound the expression as

### v3

@@ Appendix B Approximate Median
where the last inequality holds since $K=\lceil 8\cdot\log_{20\gamma^{\prime}}\left(\frac{\gamma}{4}\right)\rceil$. Notice that $K>0$ because $20\gamma^{\prime}\leq\frac{20}{27}<1$. Thus, the expected number of nodes that do not output a correct approximate median is at most $\frac{\gamma^{4}n}{256}$.
@@
Appendix C Shifting Quantiles
In this section, we prove the theoretical runtime guarantees of Theorem 2 for our robust quantile-shifting algorithm. See 2
Note that the assumptions $\beta\leq\frac{\varepsilon^{2.5}}{16}$ and $\varepsilon\leq\frac{1}{6}$ together imply that $\beta\leq\frac{\varepsilon\cdot\varepsilon^{3/2}}{16}\leq\frac{\varepsilon}{16}\cdot\frac{1}{6\sqrt{6}}=\frac{\varepsilon}{96\sqrt{6}}\leq\frac{\varepsilon}{200}$. We begin with a deterministic runtime bound for Algorithm 2.
@@ Appendix C Shifting Quantiles
Note that lines $8,9$ in Algorithm 2 are executed with probability $1$ during the first t1t-1 iterations. Indeed, $\delta$ is set to $1$ if and only if $h^{\prime}_{i}-h^{\prime}_{i+1}\leq h^{\prime}_{i}-T$ or equivalently $T\leq h^{\prime}_{i+1}$, which is true up to i=t1i=t-1 by definition of $t$. Recall the following recursive definitions:
$h_{0}^{\prime}\coloneqq 1-(\phi+\varepsilon),\qquad h_{i+1}^{\prime}\coloneqq(h_{i}^{\prime}-\beta)^{2},$
$h_{0}\coloneqq 1-(\phi+\varepsilon),\qquad h_{i+1}\coloneqq(h_{i}+\beta)^{2}.$
We now prove that $h_{t},h^{\prime}_{t}$ are close to each other, which we will later use to show that $H_{t}$ is concentrated around its expectation.
@@
For the upper bound on $h_{t-1}$, notice that by Claim 22, $h_{t-1}\geq h^{\prime}_{t-1}\geq 9/32$. Now, we will show that for all $0\leq i\leq t-2$, we have $h^{\prime}_{i}>10/19$. Suppose that by contradiction $h^{\prime}_{i}\leq 10/19$ for some such $i$. This must mean that $h^{\prime}_{i+1}\coloneqq\left(h^{\prime}_{i}-\beta\right)^{2}\leq\left(\frac{10}{19}\right)^{2}<\frac{9}{32}$, which is a contradiction. Thus, we have $h_{i}\geq h^{\prime}_{i}\geq 10/19$ for all $0\leq i\leq t-2$. Using this and applying the bounds on $h_{t-1}$ and $h^{\prime}_{t-1}$ iteratively, we get
@@
For iteration $1\leq i\leq t-1$, $\left(\frac{|H_{i}|}{n}-\beta\right)^{2}\leq\mathbb{E}\left[\frac{|H_{i+1}|}{n}\mid H_{i}\right]\leq\left(\frac{|H_{i}|}{n}+\beta\right)^{2}$
@@
Suppose that $(1-\varepsilon^{\prime\prime})h^{\prime}_{t-1}\leq\frac{|H_{t-1}|}{n}\leq(1+\varepsilon^{\prime\prime})h_{t-1}$ for some $16\beta\leq\varepsilon^{\prime\prime}\leq 1/9$. Then we have $T-\frac{3}{2}\cdot\varepsilon^{\prime\prime}\leq\mathbb{E}\left[\frac{|H_{t}|}{n}\mid H_{t-1}\right]\leq T+\frac{3}{2}\cdot\varepsilon^{\prime\prime}+\frac{1}{2}\cdot\varepsilon$.
@@
In particular $\mathbb{E}\left[\frac{|H_{t}|}{n}\mid H_{t-1}\right]=\Theta(1)$ (i.e. $\mathbb{E}\left[\frac{|H_{t}|}{n}\mid H_{t-1}\right]$ is bounded away from 0). Thus, by using Chernoff bounds with $\varepsilon^{\prime}$, we get w.h.p.
@@
In the rest of this section, we aim to show that $M_{i}$ does not change much over the course of the algorithm. First, we investigate the change in the value of $|M_{i}|$ for the first t1t-1 iterations of the algorithm.
@@
For $0\leq i<t-1$, suppose that $\frac{|H_{i}|}{n}\geq\frac{1}{2}$ and $\frac{|M_{i}|}{n}\geq\frac{|M_{0}|}{n}$, then $\mathbb{E}\left[\frac{|M_{i+1}|}{n}\mid M_{i},H_{i}\right]\geq\frac{|M_{i}|}{n}\cdot\left(1+\frac{7\varepsilon}{4}\right)$.
@@
For $0\leq i\leq t-1$, suppose that $\frac{|M_{i}|}{n}\geq\frac{|M_{0}|}{n}$, then
@@
First, note that for $0\leq i\leq t-2$, w.h.p., we must have $\frac{|H_{i}|}{n}>1/2$. Otherwise, by Lemma 24 and using Chernoff bounds with error $\varepsilon^{\prime}=1/100$, and the inequalities $\varepsilon\leq\frac{1}{6},\beta\leq\frac{\varepsilon}{200}\leq\frac{1}{100}$, we have w.h.p.
@@
Next, we show by induction that w.h.p. $\frac{|M_{i}|}{n}\geq\frac{|M_{0}|}{n}$ for $0\leq i\leq t-1$. The claim is trivially true for $i=0$. Otherwise, suppose that $\frac{|M_{i}|}{n}\geq\frac{|M_{0}|}{n}$ is true for some $i$. Using Lemma 28 and 30 with $\varepsilon^{\prime}\coloneqq\varepsilon/4$, we have
@@
By Lemmas 27 and 31, we have $\frac{1}{2}-\frac{15\varepsilon}{8}\leq\frac{|H_{t}|}{n}\leq\frac{1}{2}-\frac{\varepsilon}{8}$ and $\frac{|M_{t}|}{n}\geq 2\varepsilon$ w.h.p. Consequently, we must have $\frac{|H_{t}|}{n}+\frac{|M_{t}|}{n}\geq\frac{1}{2}+\frac{\varepsilon}{8}$. Thus, we conclude that any $\phi^{\prime}$-quantile where $\phi^{\prime}\in\left[\frac{1}{2}-\frac{\varepsilon}{8},\frac{1}{2}+\frac{\varepsilon}{8}\right]$ must be in $M_{t}$. ∎
Appendix D Approximate Mean
In this section, we prove the runtime and correctness guarantees for our adversarially robust approximate mean algorithm. We start by recalling Theorem 3.
@@ Appendix D Approximate Mean
We remark that by assumption $\beta\leq\left(\frac{\varepsilon}{100}\right)^{2.5}$ and $\delta<<\left(\frac{\varepsilon}{100}\right)^{2.5}$, which implies that $\eta\leq\left(\frac{\varepsilon}{100}\right)^{2.5}$, where $\delta$ is defined on line $1$ of Algorithm 3. Furthermore, we can rewrite $\eta$ as
@@
Recall the definitions of $\psi(t)\coloneqq\sum_{u\in V}x_{u}(t)$ and $\Phi(t)\coloneqq\sum_{u\neq v\in V}(x_{u}(t)-x_{v}(t))^{2}$. We first give two alternate expressions for $\Phi(t)$ in terms of $\psi(t)$.
@@
Now, notice that $\Phi(t+1)$ is a function of $2n$ independent random variables that correspond to the $2$ nodes chosen by each of the $n$ nodes in this round. Each of these variables appear in exactly n1n-1 of the summation terms and can affect this term by at most $M^{2}$. Thus, the effect of each of these variables on $\Phi(t+1)$ is bounded by $nM^{2}$, so the sum of squares of the effects of all these random variables is bounded by $2n^{3}M^{4}$.
@@
Appendix E Lower Bounds on Round Complexity
This section provides the proof of Proposition 5.
See 5